A cloth manufacturer produces two types of cloth, A and B.

His profit function is: Profit = 128A -4A2 + 8AB +64B -28 – 8B2

What should be the level of output of A and B to maximize profit?
Answer #1

Profit would be maximized where dP/dA = dP/dB = 0

P = 128A - 4A2 + 8AB + 64B -28 - 8B2

dP/dA = 128 - 8A + 8B = 0

dP/dB = 8A - 16B + 64 = 0

solving the above two equations simultaneously:

A = 40

B = 24